OpenStudy (anonymous):

Let f(x) = 3x+1 and g(x) = x2 Find g(f(x))

OpenStudy (anonymous):

@JA1

OpenStudy (anonymous):

i am comletely lost on this

OpenStudy (ja1):

Ugh I hate this type of math.

OpenStudy (ja1):

@joemath314159 @dumbcow

OpenStudy (anonymous):

ok thanks for looking

OpenStudy (ja1):

no problem :)

OpenStudy (anonymous):

I guess you gave up on what I was trying to show you?

OpenStudy (anonymous):

i couldnt follow you sorry @Paynesdad

OpenStudy (anonymous):

I think you have to start off understanding what f(x) means...

OpenStudy (anonymous):

f is not a variable ... it is the name of a function...A function is something that is done to a number. So when we write f(x) we are not saying f times x because f is not a variable or a number but a function.

OpenStudy (anonymous):

i have nevere been good at this im sorry for bothering you so much

OpenStudy (anonymous):

What of f(x) means is f is a function and a number will be put into that function ... the number is represented by x ... x is the input.

OpenStudy (anonymous):

@countrycuties16 don't worry about it... we all have things we struggle at...I am glad someone was there to help me improve at the things I struggled with.

OpenStudy (anonymous):

LOL It is a true fact that Einstein truly struggled with math early in his life... he turned out to be pretty good at ... :-)

OpenStudy (anonymous):

so i know that f(x)=f(3) and g(x)=x^2

OpenStudy (anonymous):

so anyway back to f(x) so we can say that when I write f(x) =3x+1 we are really saying that f is a function that take a number multiplies it by 3 and then adds 1 to it. The x is meant to represent that unknown number that we will put into f.

OpenStudy (anonymous):

Wait f(x) = f(3)....f(x) means the function f with x being put into it. and f(3) means the function f with the number 3 being put into it....

OpenStudy (anonymous):



OpenStudy (anonymous):

oh ok. [f(x)]^2 is where i got confused at before

OpenStudy (anonymous):

So since f(x) = 3x+1 then f(3) = 3(3) +1 = 10 so f(3) = 10

OpenStudy (anonymous):

Ok but lets take it one step at a time.

OpenStudy (anonymous):

ok i understand that

OpenStudy (anonymous):

So what is f(8) then ?

OpenStudy (anonymous):

f(8)=69?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

oops o messed up i multiplied 8*8=64 then 64+1=65and f(8)=65

OpenStudy (anonymous):

BUt remember ... f(x) = 3x+1 not 8x+1

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

but where did the 8 come from

OpenStudy (anonymous):

so f(8) = 3(8)+1 = 25

OpenStudy (anonymous):

so what would f(10) = ? show the work...

OpenStudy (anonymous):

3*10=30+1=31

OpenStudy (anonymous):

There you go ... so now here is a tricky one ... what is f(t)... hint the answer is not a number it is an expression with the letter t in it... Just show the same steps as before.

OpenStudy (anonymous):

3(t)=3+t=3t+1

OpenStudy (anonymous):

It would not be 3(t) but f(t) = ... but your answer is right... it should look like this...f(t) = 3t +1 ... Nowq try this f(q)

OpenStudy (anonymous):

f(q)=3(q)=3q+1 f((q)=3q+1

OpenStudy (anonymous):

Ok Don't write f(q) = 3(q) because when you say that then you are saying f(q) is 3 times q and it isn't... it is 3(q) +1 so don't write the inbetween step f(q) = 3(q) just write f(q) = 3(q) +1... does that make sense?

OpenStudy (anonymous):

ok i understand that can i finidh my other problem real quick and come back to this one because the other shouldnt take me to long to do because i undrstand how to do it